OpenStudy (anonymous):

Find the absolute value of |–1 + 3i|. I think the answer is 3, or SQRT 10. Do you agree?

myininaya (myininaya):

\[\sqrt{(-1)^2+(3)^2}\]

myininaya (myininaya):

\[|a+bi|=\sqrt{a^2+b^2}\]

OpenStudy (anonymous):

So I was correct. Thank you.